OpenStudy (sammieell):

Does anyone have the answers to this test will give a medal 1. What are the first four terms of the sequence given by the formula below? an = –2n + 4 (1 point) –6, –8, –10, –12 –2, 0, 2, 4 2, 0, –2, –4 6, 8, 10, 12 2. What is the 8th term of the sequence given by an = 3n – 5? (1 point) –2 19 24 29 3. Write a recursive definition for the sequence below. –1, 4, 9, 14, … (1 point) a1 = –1; an = an − 1 + 5 a1 = –1; an = an – 1 + 15 a1 = 14; an = an − 1 – 5 a1 = 14; an = an – 1 + 15 4. Write an explicit formula for the sequence 2, 8, 14, 20, 26, … (1 point) an = 2n – 2 an = 2n + 2 an = 4n + 2 an = 6n – 4

OpenStudy (sammieell):

Amelia is starting a new exercise routine. On the first day, she will do 8 sit-ups. On the second day, she will do 12 sit-ups. On the third day, she will do 16 sit-ups. On the fourth day, she will do 20 sit-ups. Write an explicit formula for the sequence of the number of sit-ups Amelia will do an = 3n + 2 an = 4n + 4 an = 6n – 4 an = 10n – 2 6. What is the 15th term of the sequence 20, 16, 12, 8, 4, …? (1 point) –48 –40 –36 –32 7. What is the missing term of the arithmetic sequence? 6, __, 30, … (1 point) 12 18 22 15 8. Is the sequence geometric? If so, identify the common ratio. 256, 64, 16, 4… (1 point) yes; 4 yes; yes; no 9. What is the 12th term of the sequence 3, 6, 12, 24…? (1 point) 2,048 3,072 6,144 12,288 10. What are the possible values of the missing term in the geometric sequence? 4, __, 9, … (1 point) ±5 ±6 ±13 ±36 What is the sum of the finite arithmetic series? 4 + 8 + 12 + 16 + … + 76 (1 point) 760 780 800 860 12. What is the summation notation for the series? 200 + 175 + 150 + 125 +…+(–300) (1 point) 13. What is the sum of the arithmetic series ? (1 point) 14 35 45 54 14. What is the sum of the geometric series, rounded to the nearest whole number? (1 point) 2 3 4 6 15. Does the infinite series below converge or diverge? If it converges, what is the sum? 10 + 2 + + +... (1 point) It converges; the sum is 12.5. It converges; the sum is 15. It converges; the sum is 25. It diverges